the sum of one half of the first even integer, one third of the third even integer and one fourth of the fifth even integer is equal to 12. What are the 5 consecutive even integers? : the sum of one half of the first even integer, one third of the third even integer and one fourth of the fifth even integer is equal to 12. What are the 5 consecutive even integers?
Answer by scott8148(2759)

the difference between consecutive even (or odd) integers is two __ let x=1st integer

"one half of the first even integer" __ x/2

"one third of the third even integer" __ (x+4)/3

"one fourth of the fifth even integer" __ (x+8)/4

"the sum...is equal to 12" __ x/2+(x+4)/3+(x+8)/4=12

multiplying by 12 (LCM) __ 6x+4(x+4)+3(x+8)=144 __ distributing __ 6x+4x+16+3x+24=144 __ 13x=104

dividing by 13 __ x=8 __ so the numbers are 8, 10, 12, 14, 16